# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 5. A derived Milnor-Moore theorem
$\mathrm{colim}_{\mathsf{f}\in{({\mathrm{{\mathcal{F}}in}}^{\ell+2}_{\mathrm{d}})}_{\mathrm{ndeg}}}(\tau_{>{\mathrm{n}}}({\mathcal{O}})_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{1}}\tau_{>{\mathrm{n}}}({\mathcal{O}})_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{2}}{\mathcal{O}}_{\mathsf{f}_{2}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{\ell}}{\mathcal{O}}_{\mathsf{f}_{\ell}^{-1}({\mathrm{i}})}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{\ell+1}}{\mathrm{triv}}_{\mathsf{f}_{\ell+1}^{-1}({\mathrm{i}})}),$
where ${({\mathrm{{\mathcal{F}}in}}^{\ell+2}_{\mathrm{d}})}_{\mathrm{ndeg}}\subset{\mathrm{{\mathcal{F}}in}}^{\ell+2}_{\mathrm{d}}$ is the full subcategory of sequences of maps of finite sets $\mathsf{f}:{\mathrm{d}}\xrightarrow{\mathsf{f}_{\ell+1}}{\mathrm{I}}_{\ell+1}\xrightarrow{\mathsf{f}_{\ell}}...\xrightarrow{\mathsf{f}_{1}}{\mathrm{I}}_{1}$ of length $\ell+1$ such that no map in the sequence is a bijection.
@@
$\mathrm{colim}_{\mathsf{f}\in{({\mathrm{{\mathcal{F}}in}}^{\ell+2}_{\mathrm{d}})}_{\mathrm{ndeg}}^{\prime}}({\mathrm{Lie}}_{{\mathrm{I}}_{1}}[1-|{\mathrm{I}}_{1}|]\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{1}}{\mathrm{Lie}}_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}[1-|\mathsf{f}_{1}^{-1}({\mathrm{i}})|]\otimes...$
$\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{\ell}}{\mathrm{Lie}}_{\mathsf{f}_{\ell}^{-1}({\mathrm{i}})}[1-|\mathsf{f}_{\ell}^{-1}({\mathrm{i}})|]\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{\ell+1}}{\mathrm{triv}}_{\mathsf{f}_{\ell+1}^{-1}({\mathrm{i}})}),$
where ${({\mathrm{{\mathcal{F}}in}}^{\ell+2}_{\mathrm{d}})}_{\mathrm{ndeg}}^{\prime}\subset{\mathrm{{\mathcal{F}}in}}^{\ell+2}_{\mathrm{d}}$ is the full subcategory of sequences of maps of finite sets $\mathsf{f}:{\mathrm{d}}\xrightarrow{\mathsf{f}_{\ell+1}}{\mathrm{I}}_{\ell+1}\xrightarrow{\mathsf{f}_{\ell}}...\xrightarrow{\mathsf{f}_{1}}{\mathrm{I}}_{1}$ of length $\ell+1$ such that all maps in the sequence are surjections, no map in the sequence is a bijection and the cardinality $|{\mathrm{I}}_{1}|$ of ${\mathrm{I}}_{1}$ is larger than ${\mathrm{n}}.$
For every $\mathsf{f}\in{({\mathrm{{\mathcal{F}}in}}^{\ell+2}_{\mathrm{d}})}_{\mathrm{ndeg}}^{\prime}$ and $1\leq{\mathrm{j}}\leq\ell$ the object $\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{{\mathrm{j}}}}{\mathrm{Lie}}_{\mathsf{f}_{\mathrm{j}}^{-1}({\mathrm{i}})}[1-|\mathsf{f}_{\mathrm{j}}^{-1}({\mathrm{i}})|]$ has rational homology concentrated in negative degrees. Therefore
${\mathrm{Lie}}_{{\mathrm{I}}_{1}}[1-|{\mathrm{I}}_{1}|]\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{1}}{\mathrm{Lie}}_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}[1-|\mathsf{f}_{1}^{-1}({\mathrm{i}})|]\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{\ell}}{\mathrm{Lie}}_{\mathsf{f}_{\ell}^{-1}({\mathrm{i}})}[1-|\mathsf{f}_{\ell}^{-1}({\mathrm{i}})|]\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{\ell+1}}{\mathrm{triv}}_{\mathsf{f}_{\ell+1}^{-1}({\mathrm{i}})}$
has rational homology concentrated in degrees $\leq-{\mathrm{n}}-\ell.$ Thus the cofiber ${\mathrm{C}}_{\ell}$ has rational homology concentrated in degrees $\leq-{\mathrm{n}}.$
Hence by induction on $\ell$ the rational homology of ${\mathrm{D}}_{\ell}$ is concentrated in degrees $\leq-{\mathrm{n}},$ where the case $\ell=1$ follows from the fact that ${\mathcal{D}}_{0}\simeq\tau_{>{\mathrm{n}}}({\mathcal{O}})$ has rational homology concentrated in degrees $\leq-{\mathrm{n}}.$ So $(\tau_{>{\mathrm{n}}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})_{{\mathrm{k}}}\simeq\mathrm{colim}_{\ell\geq 1}{\mathrm{D}}_{\ell}$ has rational homology concentrated in degrees $\leq-{\mathrm{n}}$.
@@
For every $[{\mathrm{n}}]\in\Delta$ let $\Delta_{\leq{\mathrm{n}}}\subset\Delta$ be the full subcategory of objects $[{\mathrm{r}}]$ with ${\mathrm{r}}\leq{\mathrm{n}}$ and $\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{\mathrm{n}}:=\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})_{|\Delta_{\leq{\mathrm{n}}}^{\mathrm{op}}}$ the restriction. There is a filtered diagram
${\mathrm{X}}\circ{\mathrm{Y}}\simeq\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{0})\to...\to\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{\mathrm{n}})\xrightarrow{\alpha_{\mathrm{n}}}$
$\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{{\mathrm{n}}+1})\to...\to\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}}))\simeq{\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{Y}},$
@@
Let ${\mathrm{L}}_{\mathrm{n}}$ be the ${\mathrm{n}}$-th latching object of $\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})$ defined as the colimit of the restriction of the functor $(\Delta^{\mathrm{op}})_{/[{\mathrm{n}}]}\to\Delta^{\mathrm{op}}\xrightarrow{\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})}{\mathcal{C}}$ to the full subcategory of surjective maps $[{\mathrm{n}}]\to[{\mathrm{k}}]$ in $\Delta$ with ${\mathrm{k}}\neq{\mathrm{n}}.$ By [23, Remark 1.2.4.3.] there is a canonical equivalence
${\mathrm{X}}\circ{\mathcal{O}}^{\circ{\mathrm{n}}}\circ{\mathrm{Y}}\simeq\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})_{\mathrm{n}}\simeq{\mathrm{L}}_{\mathrm{n}}\oplus{\mathrm{cofib}}(\alpha^{{\mathrm{n}}-1})[-{\mathrm{n}}].$
For any ${\mathrm{d}}\in\Sigma$ let ${({\mathrm{{\mathcal{F}}in}}^{{\mathrm{n}}+2}_{\mathrm{d}})}_{\mathrm{deg}},{({\mathrm{{\mathcal{F}}in}}^{{\mathrm{n}}+2}_{\mathrm{d}})}_{\mathrm{ndeg}}\subset{({\mathrm{{\mathcal{F}}in}}^{{\mathrm{n}}+2}_{\mathrm{d}})}$ be the full subcategory of sequences of maps of finite sets $\mathsf{f}:{\mathrm{d}}\xrightarrow{\mathsf{f}_{{\mathrm{n}}+1}}{\mathrm{I}}_{{\mathrm{n}}+1}\xrightarrow{\mathsf{f}_{{\mathrm{n}}}}...\xrightarrow{\mathsf{f}_{1}}{\mathrm{I}}_{1}$ of length ${\mathrm{n}}+1$ such that at least one of the maps is a bijection, no map in the sequence is a bijection, respectively.
@@
$({\mathrm{X}}\circ{\mathcal{O}}^{\circ{\mathrm{n}}}\circ{\mathrm{Y}})_{\mathrm{d}}\simeq\mathrm{colim}_{\mathsf{f}\in{\mathrm{{\mathcal{F}}in}}^{{\mathrm{n}}+2}_{\mathrm{d}}}({\mathrm{X}}_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{1}}{\mathrm{X}}_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{{\mathrm{n}}}}{\mathcal{O}}_{\mathsf{f}_{{\mathrm{n}}}^{-1}({\mathrm{i}})}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{{\mathrm{n}}+1}}{\mathrm{Y}}_{\mathsf{f}_{{\mathrm{n}}+1}^{-1}({\mathrm{i}})})$
of Remark 2.7 the summand $({\mathrm{L}}_{\mathrm{n}})_{\mathrm{d}}$ corresponds to the summand
$\mathrm{colim}_{\mathsf{f}\in{({\mathrm{{\mathcal{F}}in}}^{{\mathrm{n}}+2}_{\mathrm{d}})}_{\mathrm{deg}}}({\mathrm{X}}_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{1}}{\mathrm{X}}_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{{\mathrm{n}}}}{\mathcal{O}}_{\mathsf{f}_{{\mathrm{n}}}^{-1}({\mathrm{i}})}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{{\mathrm{n}}+1}}{\mathrm{Y}}_{\mathsf{f}_{{\mathrm{n}}+1}^{-1}({\mathrm{i}})}),$
while the summand ${\mathrm{cofib}}(\alpha^{{\mathrm{n}}-1}_{\mathrm{d}})[-{\mathrm{n}}]$ corresponds to the summand
$\mathrm{colim}_{\mathsf{f}\in{({\mathrm{{\mathcal{F}}in}}^{{\mathrm{n}}+2}_{\mathrm{d}})}_{\mathrm{ndeg}}}({\mathrm{X}}_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{1}}{\mathrm{X}}_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{{\mathrm{n}}}}{\mathcal{O}}_{\mathsf{f}_{{\mathrm{n}}}^{-1}({\mathrm{i}})}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{I}}_{{\mathrm{n}}+1}}{\mathrm{Y}}_{\mathsf{f}_{{\mathrm{n}}+1}^{-1}({\mathrm{i}})}).$
@@
By definition the functor ${\mathcal{U}}$ factors as
$\mathrm{Alg}_{{\mathrm{Lie}}}({\mathcal{C}})\simeq\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})\simeq{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\xrightarrow{{\mathrm{Grp}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}(-))}{\mathrm{Hopf}}({\mathcal{C}}),$
where the last functor is left adjoint to
${\mathrm{Grp}}({\mathrm{triv}}*_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}(-)):{\mathrm{Hopf}}({\mathcal{C}})\to{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})).$
So it is enough to see that for every ${\mathrm{X}}\in{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))$ the unit
${\mathrm{X}}\to{\mathrm{Grp}}({\mathrm{triv}}*_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}})({\mathrm{Grp}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)})({\mathrm{X}}))$
is an equivalence. Let ${\mathrm{V}}:{\mathrm{Grp}}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}}))\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathcal{C}})$ be the forgetful functor. The unit lies over the unit
${\mathrm{V}}({\mathrm{X}})\to{\mathrm{triv}}*_{{\mathrm{coComm}}_{\mathcal{C}}^{\mathrm{nu}}}({\mathrm{triv}}\circ_{{\mathrm{Lie}}(1)}({\mathrm{V}}({\mathrm{X}}))).$
The ${\mathrm{Lie}}(1)$-algebra ${\mathrm{V}}({\mathrm{X}})$ is loops on a ${\mathrm{Lie}}(1)$-algebra and so trivial by Proposition 5.17. So the claim follows from Proposition 4.13 and Lemma 5.23.
@@
In [11, Theorem 4.4.6.] Gaitsgory-Rozenblyum give a proof of Theorem 5.25 that relies on a filtration of the primitive elements.
@@
[…]
By [12, Proposition 2.8.] the $\infty$-category ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ is additive. By Remark 5.6 this implies that the forgetful functor ${\mathrm{Grp}}({\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}))\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ is an equivalence. The latter functor identifies with the forgetful functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Grp}}({\mathcal{D}}))\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}).$ So the functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\Omega):{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ is an equivalence. By Definition [23, 1.4.2.] the $\infty$-category $\mathrm{Sp}({\mathcal{D}})$ is the limit of the tower
@@
By Lemma 5.28 the top functor is an equivalence. By [23, Theorem 7.3.4.13.] the left vertical functor is an equivalence. Hence the right vertical functor $\alpha$ is an equivalence. The forgetful functor $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ is the pullback of the forgetful functor $\mathrm{Alg}_{\mathcal{O}}({\mathcal{D}})\to{\mathcal{D}}$. Hence the forgetful functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}))\to{\mathcal{C}}$ is the pullback of the forgetful functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{\mathcal{O}}({\mathcal{D}}))\to{\mathcal{D}}$ and so an equivalence.
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category. By [23, Proposition 3.2.4.7.] the object-wise symmetric monoidal structure on ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ is cartesian. Therefore the following definition is reasonable:
Notation 5.31.

### v2

@@ 4. A derived Milnor-Moore theorem
$\mathrm{colim}_{{\mathrm{f}}\in{({\mathrm{Fin}}^{\ell+2}_{d})}_{\mathrm{ndeg}}}(\tau_{>{n}}({\mathcal{O}})_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{1}}\tau_{>{n}}({\mathcal{O}})_{{\mathrm{f}}_{1}^{-1}({i})}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{2}}{\mathcal{O}}_{{\mathrm{f}}_{2}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{\mathrm{I}}_{\ell}}{\mathcal{O}}_{{\mathrm{f}}_{\ell}^{-1}({i})}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{\ell+1}}{\mathrm{triv}}_{{\mathrm{f}}_{\ell+1}^{-1}({i})}),$
where ${({\mathrm{Fin}}^{\ell+2}_{d})}_{\mathrm{ndeg}}\subset{\mathrm{Fin}}^{\ell+2}_{d}$ is the full subcategory of sequences of maps of finite sets ${\mathrm{f}}:{d}\xrightarrow{{\mathrm{f}}_{\ell+1}}{\mathrm{I}}_{\ell+1}\xrightarrow{{\mathrm{f}}_{\ell}}...\xrightarrow{{\mathrm{f}}_{1}}{\mathrm{I}}_{1}$ of length $\ell+1$ such that no map in the sequence is a bijection.
@@
$\mathrm{colim}_{{\mathrm{f}}\in{({\mathrm{Fin}}^{\ell+2}_{d})}_{\mathrm{ndeg}}^{\prime}}({\mathrm{Lie}}_{{\mathrm{I}}_{1}}[1-|{\mathrm{I}}_{1}|]\otimes\bigotimes_{{i}\in{\mathrm{I}}_{1}}{\mathrm{Lie}}_{{\mathrm{f}}_{1}^{-1}({i})}[1-|{\mathrm{f}}_{1}^{-1}({i})|]\otimes$...
$\otimes\bigotimes_{{i}\in{\mathrm{I}}_{\ell}}{\mathrm{Lie}}_{{\mathrm{f}}_{\ell}^{-1}({i})}[1-|{\mathrm{f}}_{\ell}^{-1}({i})|]\otimes\bigotimes_{{i}\in{\mathrm{I}}_{\ell+1}}{\mathrm{triv}}_{{\mathrm{f}}_{\ell+1}^{-1}({i})}),$
where ${({\mathrm{Fin}}^{\ell+2}_{d})}_{\mathrm{ndeg}}^{\prime}\subset{\mathrm{Fin}}^{\ell+2}_{d}$ is the full subcategory of sequences of maps of finite sets ${\mathrm{f}}:{d}\xrightarrow{{\mathrm{f}}_{\ell+1}}{\mathrm{I}}_{\ell+1}\xrightarrow{{\mathrm{f}}_{\ell}}...\xrightarrow{{\mathrm{f}}_{1}}{\mathrm{I}}_{1}$ of length $\ell+1$ such that all maps in the sequence are surjections, no map in the sequence is a bijection and the cardinality $|{\mathrm{I}}_{1}|$ of ${\mathrm{I}}_{1}$ is larger than n.{n}.
For every ${\mathrm{f}}\in{({\mathrm{Fin}}^{\ell+2}_{d})}_{\mathrm{ndeg}}^{\prime}$ and $1\leq{j}\leq\ell$ the object $\bigotimes_{{i}\in{\mathrm{I}}_{{j}}}{\mathrm{Lie}}_{{\mathrm{f}}_{j}^{-1}({i})}[1-|{\mathrm{f}}_{j}^{-1}({i})|]$ has rational homology concentrated in negative degrees. Therefore
${\mathrm{Lie}}_{{\mathrm{I}}_{1}}[1-|{\mathrm{I}}_{1}|]\otimes\bigotimes_{{i}\in{\mathrm{I}}_{1}}{\mathrm{Lie}}_{{\mathrm{f}}_{1}^{-1}({i})}[1-|{\mathrm{f}}_{1}^{-1}({i})|]\otimes...\otimes\bigotimes_{{i}\in{\mathrm{I}}_{\ell}}{\mathrm{Lie}}_{{\mathrm{f}}_{\ell}^{-1}({i})}[1-|{\mathrm{f}}_{\ell}^{-1}({i})|]\otimes\bigotimes_{{i}\in{\mathrm{I}}_{\ell+1}}{\mathrm{triv}}_{{\mathrm{f}}_{\ell+1}^{-1}({i})}$
has rational homology concentrated in degrees $\leq-{n}-\ell.$ Thus the cofiber $\mathrm{C}$$_{\ell}$ has rational homology concentrated in degrees $\leq-{n}.$
Hence by induction on $\ell$ the rational homology of $\mathrm{D}$$_{\ell}$ is concentrated in degrees $\leq-{n},$ where the case $\ell=1$ follows from the fact that ${\mathcal{D}}_{0}\simeq\tau_{>{n}}({\mathcal{O}})$ has rational homology concentrated in degrees $\leq-{n}.$ So $(\tau_{>{n}}({\mathcal{O}})\circ_{{\mathcal{O}}}{\mathrm{triv}})_{{k}}\simeq\mathrm{colim}_{\ell\geq 1}{\mathrm{D}}_{\ell}$ has rational homology concentrated in degrees $\leq-{n}$.
@@
For every ${n}]\in\Delta$ let $\Delta_{\leq{n}}\subset\Delta$ be the full subcategory of objects [r][{r}] with ${r}\leq{n}$ and $\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{n}:=\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})_{|\Delta_{\leq{n}}^{\mathrm{op}}}$ the restriction. There is a filtered diagram
${\mathrm{X}}\circ{\mathrm{Y}}\simeq\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{0})\to...\to\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{n})\xrightarrow{\alpha_{n}}$
$\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})^{{n}+1})\to...\to\mathrm{colim}(\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}}))\simeq{\mathrm{X}}\circ_{\mathcal{O}}{\mathrm{Y}},$
@@
Let ${\mathrm{L}}_{n}$ be the n{n}-th latching object of $\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})$ defined as the colimit of the restriction of the functor $(\Delta^{\mathrm{op}})_{/[{n}]}\to\Delta^{\mathrm{op}}\xrightarrow{\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})}{\mathcal{C}}$ to the full subcategory of surjective maps ${n}]\to[{k}]$ in $\Delta$ with ${k}\neq{n}.$ By [39, Remark 1.2.4.3.] there is a canonical equivalence
${\mathrm{X}}\circ{\mathcal{O}}^{\circ{n}}\circ{\mathrm{Y}}\simeq\mathrm{Bar}({\mathrm{X}},{\mathcal{O}},{\mathrm{Y}})_{n}\simeq{\mathrm{L}}_{n}\oplus\mathrm{cofib}(\alpha^{{n}-1})[-{n}].$
For any ${d}\in\Sigma$ let ${({\mathrm{Fin}}^{{n}+2}_{d})}_{\mathrm{deg}},{({\mathrm{Fin}}^{{n}+2}_{d})}_{\mathrm{ndeg}}\subset{({\mathrm{Fin}}^{{n}+2}_{d})}$ be the full subcategory of sequences of maps of finite sets ${\mathrm{f}}:{d}\xrightarrow{{\mathrm{f}}_{{n}+1}}{\mathrm{I}}_{{n}+1}\xrightarrow{{\mathrm{f}}_{{n}}}...\xrightarrow{{\mathrm{f}}_{1}}{\mathrm{I}}_{1}$ of length n+1{n}+1 such that at least one of the maps is a bijection, no map in the sequence is a bijection, respectively.
@@
$({\mathrm{X}}\circ{\mathcal{O}}^{\circ{n}}\circ{\mathrm{Y}})_{d}\simeq\mathrm{colim}_{{\mathrm{f}}\in{\mathrm{Fin}}^{{n}+2}_{d}}({\mathrm{X}}_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{1}}{\mathrm{X}}_{{\mathrm{f}}_{1}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{\mathrm{I}}_{{n}}}{\mathcal{O}}_{{\mathrm{f}}_{{n}}^{-1}({i})}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{{n}+1}}{\mathrm{Y}}_{{\mathrm{f}}_{{n}+1}^{-1}({i})})$
of Remark 2.7 the summand $({\mathrm{L}}_{n})_{d}$ corresponds to the summand
$\mathrm{colim}_{{\mathrm{f}}\in{({\mathrm{Fin}}^{{n}+2}_{d})}_{\mathrm{deg}}}({\mathrm{X}}_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{1}}{\mathrm{X}}_{{\mathrm{f}}_{1}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{\mathrm{I}}_{{n}}}{\mathcal{O}}_{{\mathrm{f}}_{{n}}^{-1}({i})}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{{n}+1}}{\mathrm{Y}}_{{\mathrm{f}}_{{n}+1}^{-1}({i})}),$
while the summand $\mathrm{cofib}(\alpha^{{n}-1}_{d})[-{n}]$ corresponds to the summand
$\mathrm{colim}_{{\mathrm{f}}\in{({\mathrm{Fin}}^{{n}+2}_{d})}_{\mathrm{ndeg}}}({\mathrm{X}}_{{\mathrm{I}}_{1}}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{1}}{\mathrm{X}}_{{\mathrm{f}}_{1}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{\mathrm{I}}_{{n}}}{\mathcal{O}}_{{\mathrm{f}}_{{n}}^{-1}({i})}\otimes\bigotimes_{{i}\in{\mathrm{I}}_{{n}+1}}{\mathrm{Y}}_{{\mathrm{f}}_{{n}+1}^{-1}({i})}).$
@@
The statement follows from Theorem 4.17, Corollary 4.25 and Lemma 4.26.
@@
In [18, Theorem 4.4.6.] Gaitsgory-Rozenblyum give another proof of Theorem 4.28 via a filtration of the primitive elements.
@@
By Lemma 4.31 the adjunction ${\mathrm{Barc}}:{\mathrm{Grp}}({\mathcal{D}})\rightleftarrows{\mathcal{D}}_{\ast}:{\mathrm{Cobar}}$ is an equivalence. The composition ${\mathcal{D}}_{*}\xrightarrow{{\mathrm{Cobar}}}{\mathrm{Grp}}({\mathcal{D}})\xrightarrow{\mathrm{\text{forget}}}{\mathcal{D}}_{*}$ is the loops functor $\Omega.$ So the functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\Omega):{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ factors as the equivalence ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Cobar}}):{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\simeq{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}_{*})\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Grp}}({\mathcal{D}}))$ followed by the functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Grp}}({\mathcal{D}}))\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}).$
By [19, Proposition 2.8.] the $\infty$-category ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ is additive. By Remark 4.6 this implies that the forgetful functor ${\mathrm{Grp}}({\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}))\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ is an equivalence. The latter functor identifies with the forgetful functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathrm{Grp}}({\mathcal{D}}))\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}).$ So the functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\Omega):{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})\to{\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}})$ is an equivalence. By Definition [39, 1.4.2.] the $\infty$-category ${\mathrm{Sp}}({\mathcal{D}})$ is the limit of the tower
@@
By Lemma 4.31 the top functor is an equivalence. By [39, Theorem 7.3.4.13.] the left vertical functor is an equivalence. Hence the right vertical functor $\alpha$ is an equivalence. The forgetful functor $\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\to{\mathcal{C}}$ is the pullback of the forgetful functor $\mathrm{Alg}_{\mathcal{O}}({\mathcal{D}})\to{\mathcal{D}}$. Hence the forgetful functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}}))\to{\mathcal{C}}$ is the pullback of the forgetful functor ${\mathrm{Grp}}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{\mathcal{O}}({\mathcal{D}}))\to{\mathcal{D}}$ and so an equivalence.
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category. By [39, Proposition 3.2.4.7.] the object-wise symmetric monoidal structure on $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ is cartesian. Therefore the following definition is reasonable:
Notation 4.34.